OpenStudy (anonymous):

f(X+2) where f(x)=x^2+5

OpenStudy (anonymous):

\[f(x)=x^2+5\\ f(\diamondsuit)=\diamondsuit^2+5\\ f(x+2)=(x+2)^2+5\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw